# Time-R1: Towards Comprehensive Temporal Reasoning in LLMs

### v1

@@ 3.3.1 Universal Bonuses and Penalties Design
where $\alpha$ is a decay coefficient. For Stage 1 inference, $\alpha$ is dynamically adjusted based on sample difficulty and training step (ranging between $0.07$ and $0.1$). This exponential reward structure, particularly when coupled with the dynamic $\alpha$, ensures that the reward signal clearly reflects the proximity of the inferred date to the ground truth, effectively allowing the model to perceive the magnitude of its temporal error $\Delta m(t_{p},t_{gt})$. See Section 3.3.3 and Section 5.1 for more discussion.
@@ 3.3.3 Dynamic Reward Mechanism
To address the cold-start challenge inherent in fine-tuning LLMs for specialized temporal tasks and to foster robust performance[28], particularly on more difficult examples, we employ a dynamic reward mechanism specifically during the Stage 1 RL fine-tuning process ( more discussion can be found at Section 5.1). This mechanism utilizes curriculum learning principles by adaptively adjusting the decay coefficient $\alpha$ used in the date accuracy reward component (Equation 9) based on data difficulty and training progression. This dynamic adjustment applies whenever $R_{\text{date}}$ is calculated for any Stage 1 subtask involving date inference (i.e., all four subtasks).
@@ 4.4.1 Stage 1: Foundational Temporal Reasoning Performance
Stage 1 fine-tuning effectiveness on core temporal understanding across four subtasks is presented in Table 1, showing the average total score achieved by Time-R1 ($\theta_{1}$) against baselines (see appendix C for details of fine-tuning curves for all subtasks and phases.). The results highlight the substantial benefits of our Stage 1 RL fine-tuning. Time-R1 ($\theta_{1}$) demonstrates a remarkable improvement in its overall average score, with an increase of approximately 153.0% over its base Qwen2.5-3B-Instruct model, even highly competitive to the 671B DeepSeek-R1 model with a significantly smaller size, securing the second-best performance on the challenging Event Ordering and Completion tasks. This strong performance, rivaling or exceeding much larger baselines, is largely attributed to our meticulously designed task-specific rewards. For instance, the inconsistency and diversity penalties for Event Ordering (detailed in Section 3.3.2) are pivotal. The learning curves in Appendix C also illustrate that the model’s adherence to response consistency and diversity for this task steadily improves, reflecting enhanced logical reasoning. Such effective instillation of logical mapping allows Time-R1 to compete effectively with much larger models on these complex temporal logic challenges.
Performance on the Inference task also shows considerable gains, with Time-R1 ($\theta_{1}$) clearly outperforming the ten times larger, specifically fine-tuned DeepSeek-Qwen-32B model. However, for Time-Difference estimation, despite a significant improvement of around 76.4% over the Llama-3.1-8B-Instruct model, a gap persists compared to top-tier baselines. This may be partly due to our base model’s lack of specialized mathematical reasoning pre-training, a common strength in larger, math-focused models [46, 47] (further discussed in Appendix G).
@@
Time-R1 ($\theta_{1}$, 3B) 0.6031 0.6775 0.7291 0.5857 0.4204
@@ 4.4.2 Stage 2: Future Event Time Prediction
Figure 3: Monthly Avg. Total Score $R(x,y)$ for Stage 2 Future Event Prediction (August 2024 - Feb 2025). Compares Time-R1 variants ($\theta_{2}$ and $\theta_{2}^{\prime}$) against baselines. Evaluated with $\alpha=0.1$.
The overall Stage 2 performance, measured by Average Total Score $R(x,y)$ with strict evaluation ($\alpha=0.1$), is presented in Table 2. While models show clear improvement over Stage 1 Inference tasks, likely aided by a narrower prediction time span, further significant gains prove challenging. For instance, the DS-Qwen-32B model, despite its scale and specialized complex reasoning training, scores lower than some 3B models lacking such enhancements (e.g., the base Qwen2.5-3B-Instruct), underscoring the inherent difficulty of learning extrapolation and handling post-cutoff data.Our primary model, Time-R1 ($\theta_{2}$, 3B), achieves the highest score. This strong performance, consistent across the prediction horizon (Figure 3), shows it generally outperforming most baselines, including the much larger DeepSeek-R1-671B and DeepSeek-V3-671B models. This robust result strongly supports our hypothesis that specialized, staged temporal fine-tuning enables smaller models to achieve superior performance on challenging future prediction tasks. Furthermore, these findings highlight general LLM weaknesses in temporal reasoning and underscore the efficacy and necessity of our structured training framework. The foundational understanding from Stage 1, combined with Stage 2’s predictive skill development, underpins this strong near-future temporal reasoning (see Section 5.3 for challenges facing standard LLMs). The ablation model, Time-R1-S2-Direct ($\theta_{2}^{\prime}$, 3B), also demonstrates solid performance, outperforming several baselines and indicating Stage 2 RL fine-tuning’s standalone effectiveness. See more discussion on Section 4.5.
@@
Avg. Total Score $\uparrow$ 0.6036 0.6226 0.6015 0.5997 0.7036 0.7503 0.7234 0.7697
@@ 4.4.3 Stage 3: Creative Scenario Generation Quality
Finally, we evaluate model generalization to generating plausible future scenarios—a task without explicit fine-tuning. Table 3 presents AvgMaxSim scores, quantifying the semantic plausibility of generated news scenarios against real news events (August 2024 - February 2025). Results demonstrate Time-R1 ($\theta_{2}$, 3B)’s strong generalization capability. It achieves the highest overall AvgMaxSim score, surpassing all baseline models, including the very large DeepSeek-V3-0324-671B and DeepSeek-R1-671B. Monthly scores for Time-R1 ($\theta_{2}$, 3B) also reveal consistently strong performance. This Stage 3 success, achieved without direct training on generation, underscores the S1+S2 curriculum’s effectiveness in building robust, transferable temporal reasoning. These capabilities are significant for addressing research gaps in challenging future prediction and generation tasks and demonstrate practical application value. Our ablation model, Time-R1-S2-Direct ($\theta_{2}^{\prime}$, 3B), also performs commendably, outperforming some baselines (further discussion in Section 4.5).
@@
Time-R1-S2-Direct ($\theta_{2}^{\prime}$, 3B) 47.89 46.79 49.16 47.64 48.71 48.25 48.01 46.65
Time-R1 ($\theta_{2}$, 3B) 49.22 48.10 51.19 49.97 48.86 50.16 49.29 46.95
4.5 Ablation Study on Staged Curriculum Learning
@@
The results unequivocally highlight the benefits of the full curriculum. In Future Event Time Prediction (Stage 2, Table 2, Figure 3), Time-R1 ($\theta_{2}$, 3B) (0.7697) significantly outperformed Time-R1-S2-Direct ($\theta_{2}^{\prime}$, 3B) (0.7234). This advantage persisted in Stage 3 Creative Scenario Generation (Table 3), with scores of 49.22% and 47.89% respectively. These consistent gains demonstrate that the temporal logic and event-time mapping skills instilled by Stage 1 are crucial for achieving superior predictive accuracy and generative plausibility, validating our progressive learning approach.
@@ 5 Discussion
This section delves into a detailed analysis of our proposed methodology, focusing on the efficacy of the dynamic reward curriculum learning approach, the impact of our reasoning process on response length, and the challenges standard LLMs face in advanced temporal tasks. Our findings provide empirical evidence supporting the benefits of specialized training regimes for comprehensive temporal intelligence in LLMs. Additional discussion on implementation settings (e.g., KL loss coefficients), as well as more generated examples like those shown in Figure 1, is available in Appendices D and E.
Figure 4: Impact of the Dynamic Reward Curriculum. (a) Total Score $R(x,y)$ on the Stage 1 Time-Difference Estimation task, comparing training "With Dynamic Reward" (blue, solid line) versus "Without Dynamic Reward" (red, dashed line, fixed $\alpha=0.1$). (b) Corresponding Average Response Length across all Stage 1 tasks for the same two training strategies. The y-axis for response length is broken to accommodate different scales. The dynamic reward approach not only achieves higher task scores but also promotes significantly more concise model outputs.
5.1 Dynamic Reward Curriculum Learning
Figure 4(a) empirically validates this for the Stage 1 Time-Difference Estimation task. It compares the Total Score $R(x,y)$ of our model trained with the full dynamic reward curriculum ("With Dynamic Reward") against a baseline trained with a fixed, stringent $\alpha=0.1$ from the outset ("Without Dynamic Reward"). The results clearly demonstrate the curriculum’s advantage: while the baseline model’s progress slows and converges to a lower, unstable ceiling (score $\approx 0.395$), the curriculum-trained model consistently surpasses it (around training step 70-80) and achieves a significantly higher and more stable final score (approximately 0.420). This suggests that the curriculum’s initial leniency and gradual transition to stricter evaluation criteria enable more effective exploration and learning, leading to a better mastery of the task.
@@ 5.2 Reasoning Process Matters, Not Just Response Length
Developing effective LLMs requires not only accuracy but also efficient and concise responses. Unnecessarily long outputs can signify a less refined reasoning process and increase computational overhead. Our investigation reveals that the dynamic reward curriculum, coupled with a phased learning approach (Section 3.3.3), positively influences this aspect. We employ a curriculum that progresses from simpler to more complex tasks and from lenient to stricter penalties, enabling the model to explore and establish more efficient and rational reasoning pathways.
Figure 4 offers a compelling comparison. While Figure 4(a) establishes that our dynamic reward strategy enhances task performance (higher Total Score $R(x,y)$), Figure 4(b) highlights the impact on average response length for Stage 1 tasks. A striking difference emerges: the model trained "With Dynamic Reward" (blue line, bottom panel) consistently generates significantly shorter responses throughout training compared to the model "Without Dynamic Reward" (red line, top panel, note the broken y-axis).
Specifically, the "Without Dynamic Reward" model’s responses average between approximately 240 to 320 tokens. In contrast, the "With Dynamic Reward" model’s outputs are far more concise, averaging between roughly 60 to 140 tokens, mostly remaining below 120 tokens post-initialization. This substantial reduction in length, alongside superior task performance, strongly suggests that our curriculum fosters a more efficient and focused reasoning process. The model achieves better outcomes without verbose outputs, implying a clearer, more direct approach to solving temporal tasks. Such conciseness is desirable, indicating refined understanding and potentially leading to more interpretable and computationally efficient inferences.
@@ 5.3 Challenges for Standard LLMs in Advanced Temporal Tasks
Specifically, in Stage 2 Future Event Time Prediction (Table 2, Figure 3) and Stage 3 Creative Scenario Generation (Table 3), even powerful baselines like DeepSeek-R1-671B are outperformed by our significantly smaller Time-R1 ($\theta_{2}$, 3B). For instance, Time-R1 ($\theta_{2}$, 3B) achieved a leading score of 0.7697 in Stage 2 prediction (vs. DeepSeek-R1’s 0.7503) and 49.22% in Stage 3 generation (vs. DeepSeek-V3’s 48.81%). This disparity suggests that vast knowledge, large scale, or general reasoning prowess alone do not readily translate to proficiency in predicting future event timings or creatively generating plausible future scenarios. The relatively uniform and modest performance of baselines in Stage 3, in particular, highlights a general weakness in current LLM training methodologies to effectively generalize to future-oriented generation tasks.
@@ A.2 Stage-Specific Training Configurations
  • Phase 2 (Exploration; Full Task Suite, Mixed Difficulty): Trained for 400 steps.
  • Phase 3 (Transition to Strict Evaluation; Full Task Suite): Trained for 800 steps.
@@ Appendix C Detailed Stage 1 Learning Curves and Analysis
The learning curves depicted in Figure 5 offer several key insights into the effectiveness of our methodology. Firstly, the steady increase and eventual convergence of the accuracy scores ($R_{\text{acc}}$) across all subtasks in both Phase 2 and Phase 3 underscore the benefits of our dynamic reward design and curriculum learning strategy. This carefully structured approach enables the model to progressively master complex temporal logic, gradually adapting from more lenient to stricter evaluation criteria. As noted in Section 4.4.1, this robust Stage 1 performance allows our 3B Time-R1 model to surpass numerous baseline models, many of which are ten to over two hundred times larger in parameter count (Table 1). Such strong foundational capabilities in temporal comprehension are crucial and deliberately engineered to provide a solid grounding for the subsequent, more demanding future-oriented tasks in Stage 2 (Prediction) and Stage 3 (Generation).
@@ Appendix G Limitations
Firstly, the base model for Time-R1 was not pre-trained on dedicated mathematical or logical reasoning datasets. This lack of a strong mathematical foundation may have impacted its performance on Stage 1 subtasks that involve numerical calculations, such as Time-Difference Estimation. Without this foundation, the model might be more susceptible to inconsistency penalties when predicting large time differences, potentially leading to more conservative or less accurate estimations in such scenarios.

### v2

@@ 3.3.1 Universal Bonuses and Penalties Design
where $\alpha$ is a decay coefficient. For Stage 1 inference, $\alpha$ is dynamically adjusted based on sample difficulty and training step (ranging between $0.07$ and $0.1$). This exponential reward structure, particularly when coupled with the dynamic $\alpha$, ensures that the reward signal clearly reflects the proximity of the inferred date to the ground truth, effectively allowing the model to perceive the magnitude of its temporal error $\Delta m(t_{p},t_{gt})$. See Section 3.3.3 and Section 4.5.1 for more discussion.
@@ 3.3.3 Dynamic Reward Mechanism
To address the cold-start challenge inherent in fine-tuning LLMs for specialized temporal tasks and to foster robust performance[28], particularly on more difficult examples, we employ a dynamic reward mechanism specifically during the Stage 1 RL fine-tuning process ( more discussion can be found at Section 4.5.1). This mechanism utilizes curriculum learning principles by adaptively adjusting the decay coefficient $\alpha$ used in the date accuracy reward component (Equation 9) based on data difficulty and training progression. This dynamic adjustment applies whenever $R_{\text{date}}$ is calculated for any Stage 1 subtask involving date inference (i.e., all four subtasks).
@@ 4.4.1 Stage 1: Foundational Temporal Reasoning Performance
Figure 3: Stage 1 Training Performance vs. Baselines. Training curves for Time-R1 ($\theta_{1}$) and its ablation variant, Time-R1-Fixed-Reward ($\theta_{1}^{\prime}$), evaluated against baseline models (indicated by horizontal dashed lines). Plot (a) shows the Overall Total Score across all subtasks, while plot (b) presents the Masked Time Entity Completion subtask. The solid lines demonstrate our models’ scores improving throughout the training process, ultimately surpassing the performance levels of most baseline models, including those with significantly larger scales.
The effectiveness of our Stage 1 fine-tuning on core temporal understanding is demonstrated by the training dynamics in Figure 3 (see appendix C for details of fine-tuning curves for all subtasks and phases) and the final scores in Table 1. The results highlight the substantial benefits of our Stage 1 RL fine-tuning. Time-R1 ($\theta_{1}$) demonstrates a remarkable improvement in its overall average score, with an increase of approximately 171.6% over its base Qwen2.5-3B-Instruct model.
Significantly, with these improvements, Time-R1 now outperforms the much larger DeepSeek-V3-0324-671B model and is highly competitive with the state-of-the-art 671B DeepSeek-R1 model. It secures the top performance on the demanding Completion task and the second-best performance on the challenging Event Ordering task. This strong performance, rivaling or exceeding much larger baselines, is largely attributed to our meticulously designed task-specific rewards and the dynamic reward curriculum. For instance, the inconsistency and diversity penalties for Event Ordering (detailed in Section 3.3.2) are pivotal. The learning curves in Appendix C also illustrate that the model’s adherence to response consistency and diversity for this task steadily improves, reflecting enhanced logical reasoning. Such effective instillation of logical mapping allows Time-R1 to compete effectively with much larger models on these complex temporal logic challenges.
To validate the contribution of our reward design, we include an ablation model, Time-R1-Fixed-Reward ($\theta_{1}^{\prime}$), which was trained using a static, strict reward function. As shown in Figure 3, the full Time-R1 model consistently outperforms this ablation variant, underscoring the importance of the dynamic curriculum, which will be analyzed further in Section 4.5.1.
@@
Time-R1-Fixed-Reward ($\theta_{1}^{\prime}$, 3B) 0.6259 0.6623 0.6977 0.5813 0.5621
Time-R1 ($\theta_{1}$, 3B) 0.6476 0.6815 0.7555 0.5938 0.5599
@@ 4.4.2 Stage 2: Future Event Time Prediction
Figure 4: Monthly Average Total Score $R(x,y)$ for Stage 2 Future Event Prediction (August 2024 - Feb 2025). Compares Time-R1 variants ($\theta_{2}$ and $\theta_{2}^{\prime}$) against baselines. Evaluated with $\alpha=0.1$.
@@
The overall Stage 2 performance, measured by Average Total Score $R(x,y)$ with strict evaluation ($\alpha=0.1$), is presented in Table 2. While models show clear improvement over Stage 1 Inference tasks, likely aided by a narrower prediction time span, further significant gains prove challenging. For instance, the DS-Qwen-32B model, despite its scale and specialized complex reasoning training, scores lower than some 3B models lacking such enhancements (e.g., the base Qwen2.5-3B-Instruct), underscoring the inherent difficulty of learning extrapolation and handling post-cutoff data.Our primary model, Time-R1 ($\theta_{2}$, 3B), achieves the highest score. This strong performance, consistent across the prediction horizon (Figure 4), shows it generally outperforming most baselines, including the much larger DeepSeek-R1-671B and DeepSeek-V3-671B models. This robust result strongly supports our hypothesis that specialized, staged temporal fine-tuning enables smaller models to achieve superior performance on challenging future prediction tasks. Furthermore, these findings highlight general LLM weaknesses in temporal reasoning and underscore the efficacy and necessity of our structured training framework. The foundational understanding from Stage 1, combined with Stage 2’s predictive skill development, underpins this strong near-future temporal reasoning (see Section 5.2 for challenges facing standard LLMs). The ablation model, Time-R1-S2-Direct ($\theta_{2}^{\prime}$, 3B), also demonstrates solid performance, outperforming several baselines and indicating Stage 2 RL fine-tuning’s standalone effectiveness. See more discussion on Section 4.5.2.
@@
Avg. Total Score $\uparrow$ 0.6036 0.6226 0.6015 0.5997 0.7036 0.7503 0.7331 0.7780
@@ 4.4.3 Stage 3: Creative Scenario Generation Quality
Finally, we evaluate model generalization to generating plausible future scenarios—a task without explicit fine-tuning. Table 3 presents AvgMaxSim scores, quantifying the semantic plausibility of generated news scenarios against real news events (August 2024 - February 2025). Results demonstrate Time-R1 ($\theta_{2}$, 3B)’s strong generalization capability. It achieves the highest overall AvgMaxSim score, surpassing all baseline models, including the very large DeepSeek-V3-0324-671B and DeepSeek-R1-671B. Monthly scores for Time-R1 ($\theta_{2}$, 3B) also reveal consistently strong performance. This Stage 3 success, achieved without direct training on generation, underscores the S1+S2 curriculum’s effectiveness in building robust, transferable temporal reasoning. These capabilities are significant for addressing research gaps in challenging future prediction and generation tasks and demonstrate practical application value. Our ablation model, Time-R1-S2-Direct ($\theta_{2}^{\prime}$, 3B), also performs commendably, outperforming some baselines (further discussion in Section 4.5.2).
@@
Time-R1-S2-Direct ($\theta_{2}^{\prime}$, 3B) 47.93 47.89 47.11 47.95 48.29 46.05 50.69 47.52
Time-R1 ($\theta_{2}$, 3B) 48.90 47.75 48.29 49.81 48.77 49.03 50.81 47.83
4.5 Ablation Studies
4.5.1 Impact of Dynamic Reward Mechanism
The empirical results presented in Figure 3 validate this hypothesis. The advantage of the dynamic reward curriculum is evident both in the Overall Total Score across all subtasks (Figure 3a) and in the specific Masked Time Entity Completion subtask (Figure 3b). For the overall performance, the full Time-R1 model achieves a consistently higher and more stable score than the fixed-reward ablation model. This performance gap is even more pronounced in the Completion subtask, where the fixed-reward model’s progress begins to slow and plateau around a score of 0.70. In contrast, the curriculum-trained model continues to improve, achieving a significantly higher and more stable final score of over 0.75. This suggests that the curriculum’s initial leniency and gradual transition to stricter evaluation criteria enable more effective exploration and learning, preventing convergence to a sub-optimal policy and leading to a better mastery of the task.
4.5.2 Impact of Staged Curriculum Learning
@@
The results unequivocally highlight the benefits of the full curriculum. In Future Event Time Prediction (Stage 2, Table 2, Figure 4), Time-R1 ($\theta_{2}$, 3B) (0.7780) significantly outperformed Time-R1-S2-Direct ($\theta_{2}^{\prime}$, 3B) (0.7331). This advantage persisted in Stage 3 Creative Scenario Generation (Table 3), with scores of 48.90% and 47.93% respectively. These consistent gains demonstrate that the temporal logic and event-time mapping skills instilled by Stage 1 are crucial for achieving superior predictive accuracy and generative plausibility, validating our progressive learning approach.
@@ 5 Discussion
This section delves into a detailed analysis of our proposed methodology, focusing on the impact of our reasoning process on response length, and the challenges standard LLMs face in advanced temporal tasks. Our findings provide empirical evidence supporting the benefits of specialized training regimes for comprehensive temporal intelligence in LLMs. Additional discussion on implementation settings (e.g., KL loss coefficients), as well as more generated examples like those shown in Figure 1, is available in Appendices D and E.
@@ 5.1 Reasoning Process Matters, Not Just Response Length
Figure 5: Impact of Dynamic Reward on Response Length. The average response length (in tokens) across all Stage 1 tasks during training. The model trained with our full dynamic reward mechanism ("Dynamic Reward") produces consistently and significantly more concise outputs compared to the ablation model trained with a static, fixed reward ("Fixed Reward").
Developing effective LLMs requires not only accuracy but also efficient and concise responses. Unnecessarily long outputs can signify a less refined reasoning process and increase computational overhead. Our investigation reveals that our dynamic reward mechanism (Section 3.3.3) achieves both higher accuracy and greater conciseness.
A combined analysis of our models’ performance and output length provides compelling evidence for this. As established in Section 4.5.1, our dynamic reward curriculum leads to superior task performance (Figure 3). Simultaneously, Figure 5 highlights the dramatic impact on average response length. The model trained with a fixed reward produces verbose outputs, averaging approximately 250 tokens. In stark contrast, the model trained with our dynamic reward mechanism generates significantly shorter responses, stabilizing at a much more efficient length of around 130 tokens.
This substantial reduction in length, achieved alongside superior task performance, strongly suggests that our curriculum fosters a more efficient and focused reasoning process. The model learns to achieve better outcomes without verbose outputs, implying a clearer, more direct approach to solving temporal tasks. Such conciseness is highly desirable, indicating a more refined understanding and leading to more interpretable and computationally efficient inferences.
@@ 5.2 Challenges for Standard LLMs in Advanced Temporal Tasks
Specifically, in Stage 2 Future Event Time Prediction (Table 2, Figure 4) and Stage 3 Creative Scenario Generation (Table 3), even powerful baselines like DeepSeek-R1-671B are outperformed by our significantly smaller Time-R1 ($\theta_{2}$, 3B). For instance, Time-R1 ($\theta_{2}$, 3B) achieved a leading score of 0.7780 in Stage 2 prediction (vs. DeepSeek-R1’s 0.7503) and 48.90% in Stage 3 generation (vs. DeepSeek-V3’s 48.81%). This disparity suggests that vast knowledge, large scale, or general reasoning prowess alone do not readily translate to proficiency in predicting future event timings or creatively generating plausible future scenarios. The relatively uniform and modest performance of baselines in Stage 3, in particular, highlights a general weakness in current LLM training methodologies to effectively generalize to future-oriented generation tasks.
@@ A.2 Stage-Specific Training Configurations
  • Phase 2 (Exploration; Full Task Suite, Mixed Difficulty): Trained for 500 steps.
  • Phase 3 (Transition to Strict Evaluation; Full Task Suite): Trained for 1000 steps.
@@ Appendix C Detailed Stage 1 Learning Curves and Analysis
The learning curves depicted in Figure 6 offer several key insights into the effectiveness of our methodology. Firstly, the steady increase and eventual convergence of the total scores ($R(x,y))$) across all subtasks in both Phase 2 and Phase 3 underscore the benefits of our dynamic reward design and curriculum learning strategy. This carefully structured approach enables the model to progressively master complex temporal logic, gradually adapting from more lenient to stricter evaluation criteria. As noted in Section 4.4.1, this robust Stage 1 performance allows our 3B Time-R1 model to surpass numerous baseline models, many of which are ten to over two hundred times larger in parameter count (Table 1). Such strong foundational capabilities in temporal comprehension are crucial and deliberately engineered to provide a solid grounding for the subsequent, more demanding future-oriented tasks in Stage 2 (Prediction) and Stage 3 (Generation).
@@ Appendix G Limitations
Firstly, while our work introduces Time-Bench, a large-scale, open-source dataset designed to facilitate comprehensive temporal reasoning research, a potential limitation lies in the scope of evaluation. Spanning a decade of news data and comprising over 200,000 examples across multiple temporal tasks, Time-Bench provides a robust benchmark for evaluating the capabilities demonstrated by Time-R1. However, validating the effectiveness and generalization capabilities of our model on a wider array of external temporal reasoning benchmarks and diverse datasets would further strengthen our findings and provide stronger evidence for the robustness of our proposed training framework.